Fermi National Laboratory


Accelerator Update for April 11, 2005 - April 13, 2005

Monday April 11
The day shift began with Operations monitoring store 4089 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with the MiniBooNE and the Meson MCenter experiments taking beam.

Pbar experts began a study at 9:57 AM.

The Neutron Therapy Facility (NTF) requested beam at 10 AM for calibration. They were done two hours later.

A Tevatron (TeV) expert began a study at 10:03 AM that would fly a wire once every hour.

An expert took a Booster RF station (BRF5) off line at 10:50 AM to investigate and repair a LCW leak on a contactor.

Operations turned off Booster at 11:31 AM for scheduled breaker work. The work took one hour.

Operations established beam to MiniBooNE and MCenter at 12:43 PM.

Pbar experts resumed their studies at 12:49 PM.

At 1:10 PM, the TeV suffered a sector D1 quench due to a safety system trip. The Main Injector (MI) beam permit also tripped off and the stash was lost. Later, investigations found that safety technicians were working on the NuMI safety system.

Operations prepared the Collider Detector at Fermilab (CDF and D-Zero (D0) for an access at 2:07 PM. They were both out of access by 3:44 PM

Operations put the TeV into shot setup at 5:08 PM.

At 6:47 PM, after loading the antiprotons, the TeV suffered a sector E4 quench due to a sequencer problem.

MCenter resumed taking beam at 7:12 PM

TeV experts began a study period at 10:11 PM.

Tuesday April 12
The midnight shift began with TeV experts conducting studies, with Operations monitoring the stash, and with MiniBooNE and MCenter taking beam.

Operations put the TeV into shot setup at 1:15 AM.

Operations established store 4095 at 3:21 AM with an initial luminosity of 61.05E30. Pbar resumed stacking.

MCenter resumed taking beam at 3:44 AM.

At 3:44 AM, a Booster device (MP01) tripped off and wouldn't reset. Operators contacted an expert. This held off MCenter beam.

Operations got MP01 back on and resumed sending beam to MCenter at 7 AM.

MP01 tripped off at 7:48 AM. Operations contacted an expert.

Operations put the Recycler into shot setup to stash antiprotons at 7:57 AM.

At 8:25 AM, a Booster RF station (BRF16) tripped off and wouldn't reset. Operations bypassed the station.

Pbar experts reported at 9:53 AM that they were tuning the machine.

After many delays, the Recycler stash was completed at 11:28 AM.

Pbar resumed stacking at 11:46 AM.

At 4:58 PM, a Linac quadrupole power supply alarmed on an over temperature. Operators investigated and then called in an expert. The expert had the supply back on by 5:48 PM.

Operations terminated store 4095 at 10:38 PM. Operators prepared CDF and D0 for access into their collision halls.

At 11:18 PM, a Debuncher kicker (D:EKIK) stopped firing. Operators investigated and then called in and expert.

Wednesday April 13
The midnight shift began with Operations monitoring the stash and the stack, with stacking off due to a kicker problem, and with MiniBooNE and MCenter taking beam.

The kicker expert reported finding a cable short, which he repaired. Pbar resumed stacking at 1:31 AM.

Operations put the TeV into shot setup at 2:14 AM.

Operations established store 4097 at 4:29 AM with an initial luminosity of 78.64E30. Pbar resumed stacking and MCenter resumed taking beam.

The Plan for Wednesday
The plan for today is to stash the Recycler this morning and then allow Pbar and MI experts to conduct studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/13/2005   email Fermilab